OpenStudy (anonymous):

D = {x|x is a whole number} E = {x|x is a perfect square between 1 and 9} F = {x|x is an even number greater than or equal to 2 and less than 9} E ∪ F is {2, 4, 6, 8} True False

OpenStudy (displayerror):

What number(s) belong to set E? What number(s) belong to set F? The symbol between E and F represents the union. This set contains all of the numbers that are in E OR F. As an example, \[A = \left\{ 1, 2, 3 \right\}\] \[B = \left\{ 4, 5, 6 \right\}\] \[A \cup B = \left\{ 1, 2, 3, 4, 5, 6 \right\}\]

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

@johnweldon1993